OpenStudy (anonymous):

The volume of the box is 60 cubic inches. What are the length, width, and height of the box? HELP!

OpenStudy (anonymous):

Is the box a cube?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay, that makes it possible to calculate. \[l^3 = 60\] Solve for l, and you have your answer.

OpenStudy (anonymous):

ya..

OpenStudy (anonymous):

Is it 30?

OpenStudy (anonymous):

???? I'm completely feel stuck on this.

OpenStudy (anonymous):

That's okay. We can work through it together. \[l^3 = 60\] What should you do to both sides to isolate the l?

OpenStudy (anonymous):

\[l = \sqrt[3]{60}\]

OpenStudy (anonymous):

Divide.

OpenStudy (anonymous):

It is the cubic root, so we do not divide in this case. The opposite of square is root, so we take the third root of 60.

OpenStudy (anonymous):

Which would be 15 right?

OpenStudy (anonymous):

The answer should come out as around 3.9.

OpenStudy (anonymous):

You can always check your answer by putting your solution to the power of 3.

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

I see... I had to round it off.

OpenStudy (anonymous):

All you need to do is put this in your calculator: \[\sqrt[3]{60}\]

OpenStudy (anonymous):

I didn't get that on my calculator ..

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

3.27

OpenStudy (anonymous):

I'm not sure how you got that, but you know how to work out the answer, and that's the important thing :)